### v2

# On the Subpacketization Level of the Banawan–Ulukus Multi-Message PIR Scheme

Abstract
This note analyzes a linear recursion that arises in the computation of the subpacketization level for the multi-message PIR scheme of Banawan and Ulukus. We derive an explicit representation for the normalized subpacketization level $L$, whose smallest integer multiple yields the subpacketization level of the scheme, in terms of the number of servers NN, the total number of messages $K$, and the number of demand messages $D$. The resulting formula shows that $L$ is a polynomial in NN with nonnegative coefficients, and its leading term is $N^{K-D+1}/D$.
I Setup and Main Results
Fix integers $K>D>1$ and $N>1$, and let $\{L_{1},\dots,L_{K}\}$ be a sequence satisfying
$L_{K}=(N-1)^{K-D},$ (1)
$L_{K-D+1}=L_{K-D+2}=\cdots=L_{K-1}=0,$ (2)
$L_{j}=\frac{1}{N-1}\sum_{i=1}^{D}\binom{D}{i}\,L_{i+j},\quad j\in[1:K-D].$ (3)
Define
$L\triangleq\frac{N}{D}\sum_{t=1}^{K}\binom{K}{t}L_{t}\;-\;\frac{N}{D}\sum_{t=1}^{K-D}\binom{K-D}{t}L_{t}.$ (4)
The quantity $L$ in (4) is the normalized subpacketization level of the multi-message PIR scheme in [1], meaning that the scheme’s subpacketization level is the smallest integer multiple of $L$ [2]. Here NN denotes the number of servers, $K$ the total number of messages, and $D$ the number of demand messages.
Let
$T\triangleq K-D+1,\quad S\triangleq\left\lfloor\frac{T(D-1)}{D}\right\rfloor.$
For each $n\in[0:T(D-1)]$, define
$a_{n}\triangleq[x^{n}](1+x+\cdots+x^{D-1})^{T},$ (5)
where $x^{n}]p(x)$ denotes the coefficient of $x^{n}$ in the polynomial $p(x)$. Equivalently,
$(1+x+\cdots+x^{D-1})^{T}=\sum_{n=0}^{T(D-1)}a_{n}x^{n}.$
Theorem 1.
The quantity $L$ in (4) can be written as
$L=\frac{1}{D}\sum_{k=0}^{S}a_{kD}\,N^{T-k}.$ (6)
In particular, $L$ is a polynomial in NN with nonnegative coefficients, and its leading term is $N^{T}/D$.
II Proof of Theorem 1
Define the reversed sequence
$M_{t}\triangleq L_{K-t},\quad t\in[0:K-1].$
Then (1) and (2) become
$M_{0}=(N-1)^{K-D},\quad M_{1}=M_{2}=\cdots=M_{D-1}=0.$ (7)
Lemma 1.
For each $t\in[D:K-1]$, the recursion (3) implies
$NM_{t}=\sum_{i=0}^{D}\binom{D}{i}M_{t-i}.$ (8)
Proof.
Fix $t\in[D:K-1]$} and set j=Kt{j=K-t}. Then $j\in[1:K-D]$}, so (3) applies:
$L_{K-t}=\frac{1}{N-1}\sum_{i=1}^{D}\binom{D}{i}L_{i+K-t}.$
Since $L_{K-t}=M_{t}$ and $L_{i+K-t}=L_{K-(t-i)}=M_{t-i}$, we have
$M_{t}=\frac{1}{N-1}\sum_{i=1}^{D}\binom{D}{i}M_{t-i},$
or equivalently $(N-1)M_{t}=\sum_{i=1}^{D}\binom{D}{i}M_{t-i}$. Adding the $i=0$ term $\binom{D}{0}M_{t}=M_{t}$ to both sides gives
$NM_{t}=\sum_{i=0}^{D}\binom{D}{i}M_{t-i},$
which matches (8). ∎
To solve (8), take $M_{t}=r^{t}$ with $r\neq 0$. Substitution gives
$\sum_{i=0}^{D}\binom{D}{i}r^{t-i}=Nr^{t}$ $\;\Longleftrightarrow\;\sum_{i=0}^{D}\binom{D}{i}r^{-i}=N$
$\;\Longleftrightarrow\;\left(1+r^{-1}\right)^{D}=N.$
Let $s>0$ satisfy $s^{D}=N$ and, for each $m\in[0:D-1]$, define
$\omega_{m}\triangleq e^{\mathrm{i}2\pi m/D},\quad u_{m}\triangleq\omega_{m}s.$
Lemma 2.
For any integer $P$,
$\sum_{m=0}^{D-1}\omega_{m}^{P}=\begin{cases}D,&D\mid P,\\ 0,&D\nmid P.\end{cases}$
Proof.
If $D\mid P$, then $\omega_{m}^{P}=(\omega_{m}^{D})^{P/D}=1$ for every $m\in[0:D-1]$}, and hence $\sum_{m=0}^{D-1}\omega_{m}^{P}=D$. If $D\nmid P$, then $\omega_{1}^{P}\neq 1$ and, since $\omega_{m}=\omega_{1}^{m}$,
$\sum_{m=0}^{D-1}\omega_{m}^{P}=\sum_{m=0}^{D-1}(\omega_{1}^{P})^{m}=\frac{1-(\omega_{1}^{P})^{D}}{1-\omega_{1}^{P}}=\frac{1-(\omega_{1}^{D})^{P}}{1-\omega_{1}^{P}}=0,$
which proves the lemma. ∎
Lemma 3.
The solutions to $(1+r^{-1})^{D}=N$} are
$r_{m}\triangleq\frac{1}{u_{m}-1},\quad m\in[0:D-1].$ (9)
Proof.
Let $r$ be any solution. Since $N=s^{D}$, the equation $(1+r^{-1})^{D}=s^{D}$ implies $1+r^{-1}=\omega_{m}s$ for some $m\in[0:D-1]$}. Therefore,
$r=\frac{1}{\omega_{m}s-1}=\frac{1}{u_{m}-1}.$
Conversely, each $r_{m}$ in (9) satisfies $1+r_{m}^{-1}=u_{m}$ and hence $(1+r_{m}^{-1})^{D}=u_{m}^{D}=(\omega_{m}s)^{D}=s^{D}=N$. ∎
Therefore, the general solution to (8) has the form
$M_{t}=\sum_{m=0}^{D-1}c_{m}r_{m}^{t},$ (10)
where the coefficients $\{c_{m}\}_{m=0}^{D-1}$ are determined by the initial conditions (7).
Lemma 4.
With coefficients
$c_{m}\triangleq\frac{(N-1)^{K-D}}{D}\cdot\frac{(u_{m}-1)^{D-1}}{u_{m}^{D-1}},\quad m\in[0:D-1],$ (11)
the sequence $\{M_{t}\}_{t=0}^{K-1}$ defined by (10) satisfies the initial conditions (7).
Proof.
Substituting (11) into (10) and using $r_{m}=1/(u_{m}-1)$,
$M_{t}$ $=\sum_{m=0}^{D-1}\frac{(N-1)^{K-D}}{D}\cdot\frac{(u_{m}-1)^{D-1}}{u_{m}^{D-1}}\cdot\frac{1}{(u_{m}-1)^{t}}$
$=\frac{(N-1)^{K-D}}{D}\sum_{m=0}^{D-1}\frac{(u_{m}-1)^{D-1-t}}{u_{m}^{D-1}}.$
For each $t\in[0:D-1]$, expanding $(u_{m}-1)^{D-1-t}$ and dividing by $u_{m}^{D-1}$ gives
$\frac{(u_{m}-1)^{D-1-t}}{u_{m}^{D-1}}$
$\quad=\sum_{l=0}^{D-1-t}\binom{D-1-t}{l}u_{m}^{l-(D-1)}(-1)^{D-1-t-l}.$
Therefore,
$\sum_{m=0}^{D-1}\frac{(u_{m}-1)^{D-1-t}}{u_{m}^{D-1}}$
$\quad=\sum_{l=0}^{D-1-t}\binom{D-1-t}{l}(-1)^{D-1-t-l}\sum_{m=0}^{D-1}u_{m}^{l-(D-1)}.$ (12)
Since $u_{m}=\omega_{m}s$}, for any integer $P$ we have ${\sum_{m=0}^{D-1}u_{m}^{P}}={s^{P}\sum_{m=0}^{D-1}\omega_{m}^{P}}$. By Lemma 2, $\sum_{m=0}^{D-1}\omega_{m}^{P}$ equals $D$ if $D\mid P$, and equals zero otherwise.
For t=0{t=0}, the exponents l(D1){l-(D-1)} range over $\{-(D-1),\ldots,0\}$}. The only multiple of $D$ in this set is 0, which occurs at l=D1l=D-1 with coefficient 11. Thus (12) equals $D$, and hence $M_{0}=(N-1)^{K-D}$}.
For $1\leq t\leq D-1$}, the exponents l(D1){l-(D-1)} range over $\{-(D-1),\ldots,-t\}$}, none of which is divisible by $D$. Therefore (12) equals zero, and consequently $M_{t}=0$. ∎
Since $L_{t}=M_{K-t}$, we can rewrite the sums in (4) as
$\sum_{t=1}^{K}\binom{K}{t}L_{t}$ $=\sum_{t=0}^{K-1}\binom{K}{t}M_{t},$ (13)
$\sum_{t=1}^{K-D}\binom{K-D}{t}L_{t}$ $=\sum_{t=D}^{K-1}\binom{K-D}{t-D}M_{t}.$ (14)
Lemma 5.
For each $m\in[0:D-1]$},
$\sum_{t=0}^{K-1}\binom{K}{t}r_{m}^{t}$ $=(1+r_{m})^{K}-r_{m}^{K},$ (15)
$\sum_{t=D}^{K-1}\binom{K-D}{t-D}r_{m}^{t}$ $=r_{m}^{D}\left((1+r_{m})^{K-D}-r_{m}^{K-D}\right).$ (16)
Proof.
By the binomial theorem,
$\sum_{t=0}^{K}\binom{K}{t}r_{m}^{t}=(1+r_{m})^{K}.$
Dropping the $t=K$ term yields (15). For (16), factoring out $r_{m}^{D}$ and reindexing gives
$\sum_{t=D}^{K-1}\binom{K-D}{t-D}r_{m}^{t}$ $=r_{m}^{D}\sum_{t=D}^{K-1}\binom{K-D}{t-D}r_{m}^{t-D}$
$=r_{m}^{D}\sum_{t=0}^{K-D-1}\binom{K-D}{t}r_{m}^{t}$
$=r_{m}^{D}\left((1+r_{m})^{K-D}-r_{m}^{K-D}\right),$
where the last step follows by applying (15) with $K$ replaced by KDK-D. ∎
Proposition 1.
The quantity $L$ in (4) can be written as
$L=\frac{N}{D}\sum_{m=0}^{D-1}c_{m}\,(N-1)\,r_{m}^{D}(1+r_{m})^{K-D}.$ (17)
Proof.
Starting from (4) and using (13) and (14),
$L=\frac{N}{D}\left(\sum_{t=0}^{K-1}\binom{K}{t}M_{t}-\sum_{t=D}^{K-1}\binom{K-D}{t-D}M_{t}\right).$
Substituting $M_{t}=\sum_{m=0}^{D-1}c_{m}r_{m}^{t}$ and exchanging sums,
$L=\frac{N}{D}\sum_{m=0}^{D-1}c_{m}\left(\sum_{t=0}^{K-1}\binom{K}{t}r_{m}^{t}-\sum_{t=D}^{K-1}\binom{K-D}{t-D}r_{m}^{t}\right).$
Applying Lemma 5,
$\sum_{t=0}^{K-1}\binom{K}{t}r_{m}^{t}-\sum_{t=D}^{K-1}\binom{K-D}{t-D}r_{m}^{t}$
$\quad=\left((1+r_{m})^{K}-r_{m}^{K}\right)-r_{m}^{D}\left((1+r_{m})^{K-D}-r_{m}^{K-D}\right)$
$\quad=(1+r_{m})^{K-D}\left((1+r_{m})^{D}-r_{m}^{D}\right).$
By Lemma 3, $(1+r_{m}^{-1})^{D}=N$, equivalently $(1+r_{m})^{D}=Nr_{m}^{D}$, and thus $(1+r_{m})^{D}-r_{m}^{D}=Nr_{m}^{D}-r_{m}^{D}=(N-1)r_{m}^{D}$, which establishes (17). ∎
Using $r_{m}=1/(u_{m}-1)$ gives $1+r_{m}=u_{m}/(u_{m}-1)$ and hence
$r_{m}^{D}(1+r_{m})^{K-D}=\frac{u_{m}^{K-D}}{(u_{m}-1)^{K}}.$
Substituting (11) into (17) yields
$L=\frac{N(N-1)^{T}}{D^{2}}\sum_{m=0}^{D-1}\frac{u_{m}^{T-D}}{(u_{m}-1)^{T}},$ (18)
where T=KD+1T=K-D+1.
Rewriting each summand in (18), we have
$\frac{u_{m}^{T-D}}{(u_{m}-1)^{T}}=u_{m}^{-D}\bigl(1-u_{m}^{-1}\bigr)^{-T}.$
Using $u_{m}=\omega_{m}s$, we obtain $u_{m}^{-D}=(\omega_{m}s)^{-D}=s^{-D}=N^{-1}$ and $1-u_{m}^{-1}=1-\omega_{m}^{-1}s^{-1}$. Substituting into (18) gives
$L=\frac{(N-1)^{T}}{D^{2}}\sum_{m=0}^{D-1}\bigl(1-\omega_{m}^{-1}s^{-1}\bigr)^{-T}.$
Since $\{\omega_{m}^{-1}:m\in[0:D-1]\}=\{\omega_{m}:m\in[0:D-1]\}$, reindexing yields
$L=\frac{(N-1)^{T}}{D^{2}}\sum_{m=0}^{D-1}\bigl(1-\omega_{m}s^{-1}\bigr)^{-T}.$ (19)
Lemma 6.
For each $m\in[0:D-1]$,
$\frac{1-N^{-1}}{1-\omega_{m}s^{-1}}=\sum_{t=0}^{D-1}\omega_{m}^{t}s^{-t}.$
Proof.
Multiplying $(1-\omega_{m}s^{-1})$ by $\sum_{t=0}^{D-1}\omega_{m}^{t}s^{-t}$ gives
$(1-\omega_{m}s^{-1})\sum_{t=0}^{D-1}\omega_{m}^{t}s^{-t}$ $=\sum_{t=0}^{D-1}\omega_{m}^{t}s^{-t}-\sum_{t=0}^{D-1}\omega_{m}^{t+1}s^{-(t+1)}$
$=1-\omega_{m}^{D}s^{-D}=1-N^{-1},$
since $\omega_{m}^{D}=1$ and $s^{D}=N$. Dividing both sides by $(1-\omega_{m}s^{-1})$} yields the identity. ∎
For each $n\in[0:T(D-1)]$, define $a_{n}$ as in (5), so that
$(1+x+\cdots+x^{D-1})^{T}=\sum_{n=0}^{T(D-1)}a_{n}\,x^{n}.$ (20)
In particular, $a_{n}\geq 0$, since $a_{n}$ equals the number of $T$-tuples $(n_{1},\dots,n_{T})\in[0:D-1]^{T}$ such that $\sum_{l=1}^{T}n_{l}=n$.
Fix $m\in[0:D-1]$. By Lemma 6,
$\frac{(1-N^{-1})^{T}}{(1-\omega_{m}s^{-1})^{T}}=\left(\sum_{t=0}^{D-1}\omega_{m}^{t}s^{-t}\right)^{T}.$ (21)
Expanding the $T$th power using (20) gives
$\left(\sum_{t=0}^{D-1}\omega_{m}^{t}s^{-t}\right)^{T}=\sum_{n=0}^{T(D-1)}a_{n}\,\omega_{m}^{n}\,s^{-n}.$ (22)
Substituting (22) into (21) gives
$\frac{(1-N^{-1})^{T}}{(1-\omega_{m}s^{-1})^{T}}=\sum_{n=0}^{T(D-1)}a_{n}\,\omega_{m}^{n}\,s^{-n}.$ (23)
Averaging (23) over $m\in[0:D-1]$ and applying Lemma 2 yields
$\frac{1}{D}\sum_{m=0}^{D-1}\frac{(1-N^{-1})^{T}}{(1-\omega_{m}s^{-1})^{T}}$ $=\sum_{n=0}^{T(D-1)}a_{n}\,s^{-n}\left(\frac{1}{D}\sum_{m=0}^{D-1}\omega_{m}^{n}\right)$
$=\sum_{k=0}^{S}a_{kD}\,s^{-kD}$
$=\sum_{k=0}^{S}a_{kD}\,N^{-k},$ (24)
where $S=\lfloor T(D-1)/D\rfloor$.
Returning to (19) and using $(N-1)^{T}=N^{T}(1-N^{-1})^{T}$,
$L$ $=\frac{(N-1)^{T}}{D^{2}}\sum_{m=0}^{D-1}(1-\omega_{m}s^{-1})^{-T}$
$=\frac{N^{T}}{D^{2}}\sum_{m=0}^{D-1}\frac{(1-N^{-1})^{T}}{(1-\omega_{m}s^{-1})^{T}}$
$=\frac{N^{T}}{D}\left(\frac{1}{D}\sum_{m=0}^{D-1}\frac{(1-N^{-1})^{T}}{(1-\omega_{m}s^{-1})^{T}}\right).$
Substituting (24) yields
$L=\frac{N^{T}}{D}\sum_{k=0}^{S}a_{kD}\,N^{-k}=\frac{1}{D}\sum_{k=0}^{S}a_{kD}\,N^{T-k},$
which matches (6).
Since (6) is a finite sum of nonnegative integer powers of NN with coefficients $a_{kD}\geq 0$ for all $k\in[0:S]$, it follows that $L$ is a polynomial in NN with nonnegative coefficients. Moreover, the highest power of NN occurs at $k=0$, and
$a_{0}=[x^{0}](1+x+\cdots+x^{D-1})^{T}=1,$
so the leading term of $L$ is $N^{T}/D$. This completes the proof of Theorem 1.